# Private Federated Learning using Preference-Optimized Synthetic Data

### v1

@@ 5.1 Main Results
Table 1 lists the next-token prediction accuracy achieved by baseline models (DP-FedAv, DP-FTRL, Private Evolution) and POPri. In this table, we assume full participation (no client sampling) for fair comparison to baselines, some of which do not have client sampling versions. We find that POPri outperforms all the baseline algorithms. Furthermore, POPri closes the gap between fully private learning ($\epsilon=0$) and fully non-private learning ($\epsilon=\infty$) by 48-69%, depending on the setting. In contrast, PE closes the gap by between 31-49%, and PE+SFT does similarly or worse than PE. For all methods tested, the measured accuracy values do not depend strongly on $\epsilon$. This has been observed in prior work on DP synthetic data using LLMs (Xie et al., 2024; Hou et al., 2024). POPri outperforms Private Evolution (Aug-PE) even when holding our synthetic sample budget to 2000. Note that synthetic samples are cheap in POPri (we could generate many more) because we have access to the full model, while Xie et al. (2024) only assume access to a model API.
@@
For FedAvg, each round the sampled clients download and upload the downstream model, which in our case is DistilGPT2. This is an 82M (82 million) parameter model leading to a download and upload cost of 82M floats. The client runtime cost comes from local gradient computation, and server runtime is negligible because the server only needs to average model deltas from the clients. For PE, the communication cost comes from each client downloading $K=1800$ sentence embeddings of size 384 resulting in a download cost of 700K (700,000) floats, and uploading a histogram of size 1800 resulting in an upload cost of 1800 floats. The client runtime cost comes from calculating a nearest neighbors histogram and the server runtime cost for PE comes mainly from using the LLM $\Psi$ to generate synthetic samples each round. In POPri each client downloads $K\times J=1800\times 10$ sentence embeddings for a download cost of 7M floats and uploads a vector of size 18,000 for an upload cost of 18,000 floats. The client runtime cost of POPri comes from calculating the cosine similarities, and the server runtime mainly comes from both using $\Psi$ to generate synthetic samples and running DPO preference optimization.
@@ 5.2 Ablations
[…]
@@
Synthetic datasets are often generated using a language model distinct from the one being aligned (Guo et al., 2024), making the alignment phase inherently off-policy as the model evolves during training. This is reflected in the synthetic data, where the FID score (relative to a held-out evaluation set) worsens after improving. Figure 4 shows PCA visualizations of synthetic data embeddings across alignment iterations, while the left panels plot the FID score and median distance to the medoid in the PCA space. The data distribution transitions from being initially clustered to (roughly) matching the true data distribution, back to being clustered, likely due to overfitting. Early stopping based on validation metrics can help avoid this.
Preference optimization algorithms.
The POPri algorithm constructs a preference dataset out of DP scored synthetic data for model alignment. We explore two popular alignment methods: DPO and IPO. IPO (Gheshlaghi Azar et al., 2024) is proposed to address the over-fitting to the preference dataset that is common for DPO. As shown in Fig 4, over-fitting problems exist in our setting and lead to clustering of the data distributions that don’t represent the true distribution. However, empirically, we find that DPO outperforms IPO in our setting (details in Section E.4).
@@ 6 Conclusion
Private on-device learning is important when data is stored on edge devices with hardware, storage, and privacy constraints. We propose POPri, which recasts synthetic data-based approaches for private on-device learning as an LLM preference optimization problem. POPri makes several novel design choices in how it gathers and utilizes client feedback to generate DP synthetic data, which is used to finetune a downstream on-device model. POPri outperforms DP-FL and synthetic data baselines on the downstream next-word-prediction task, including on a large-scale LargeFedBench, a new federated benchmark we have curated.
@@
In this paper, we train synthetic data models satisfying differential privacy guarantees. When using differential privacy as a tool for protecting user data, it is important to communicate to users what the privacy guarantees mean to be able to obtain informed consent. The algorithms in this paper also use LLMs, which were trained on large scale public text data. While this data was public, explicit consent may not have been given for its use in training the models. The algorithms using LLMs in the paper make no claims about the privacy guarantees of data used in the pretraining of the LLMs. While our work aims to show how synthetic data can be useful for federated learning, it also poses a number of ethical risks, including the generation of biased or harmful content. In particular, our method (and all variants of private evolution) inherits the biases and undesirable aspects of the public LLM. For example, suppose the public LLM only generates text in English, but some clients’ private data is all in Spanish. In these settings, clients would be forced to vote on synthetic samples, even if potentially none of them are relevant to the client. This may cause the client to contribute data reinforcing a model that is actively not useful (or even harmful) to the client. In contrast, DP-SGD methods do not suffer from this shortcoming, because they do not rely on a public LLM. This problem raises an important point—how can we design DP synthetic data algorithms in which clients can stem the biases or failures of the public LLM, based on their own data? This important question is beyond the scope of the current paper.
@@ Appendix A Algorithmic Details
1:  Input: Embeddings of private client data $E_{i}$ for ${i\in\mathcal{S}^{t}}$, embeddings of synthetic data ${\rm E}_{syn}$, total synthetic samples $M=K\times J$ $\rm{Scores}$ $\leftarrow\boldsymbol{0}^{M}$
2:  ${\rm Scores}[j]$ = $(1/|E_{i}|)\sum_{e_{pri}\in E_{i}}\frac{\langle e_{pri},e_{j}\rangle}{\|e_{pri}\|\|e_{j}\|}$ for $e_{j}\in E_{syn}$
@@ B.1 Model and Hyperparameters
In each round we generate 18000 synthetic data samples for the clients to evaluate. This is accomplished with 1800 prompts, each generating 10 samples for clients to rank. We select the 1st and 5th ranked sample for a given prompt for the “selected” and “rejected” data samples in the DPO preference dataset. We describe the experiments regarding which rank to use for constructing the preference dataset in detail in Appendix Section E.5. To test the scaling relation with the number of clients per round and the total number of clients participating in the training, we set up the parameters and privacy budget shown in Table D.1. The ‘all-MiniLM-L6-v2’ sentence transformer model is used as the embedding model in POPri. We note that we adopt “sentence-t5-base” sentence transformer for PubMed during the step of fine-tuning ${\rm BERT}_{small}$, which follows the setting in AUG-PE. We ensure POPri follows privacy guarantee of $(\epsilon,\delta)$-DP = (1, $3\times 10^{-6}$) or (7, $3\times 10^{-6}$) for both the bioRxiv and the Congressional Speeches datasets and run with 20 iterations for DP-FedAv, DP-FTRL, PrE-Text for comparison. For AUG-PE, we set $(\epsilon,\delta)$-DP = (1, $2.72\times 10^{-6}$) or (4, $2.72\times 10^{-6}$). PubMed experiments are run with 10 iterations.
@@
  • For BioRxiv &\&& Congressional Speeches, we fine-tuned the pre-trained DistillGPT2 for next-token prediction. We set the max sequence length as 64, number of generated synthetic data as 1,000,000, the batch size as 160, the learning rate as $2e^{-4}$, and the number of epochs as 80.
[…]
@@ B.2 Prompt Design
To compare with other data generator methods, we adopt the prompts used in the baseline models against which we compare. We generate the synthetic data using an approach similar to that in PrE-Text (Hou et al., 2024). Figure 5 shows an example of the prompt we use for prompting Llama-3B for generating synthetic data. For PubMed, while running POPri, we still adopt the prompt shown in Figure 5 but reduce the number of examples to two in order to accommodate longer sequence lengths.
@@ Appendix C Implementation Details of Baseline Models
In this section we provide implementation details for the baseline algorithms. We use two DP-FL baselines: DP-FedAvg and DP-FTRL. For the PE baseline, we implement PrE-Text (Hou et al., 2024) for the evaluations on the bioRxiv and Congressional Speeches datasets. Because the PrE-Text evaluation is focused on datasets with samples with max sequence length of 64 and the PubMed dataset has samples with longer sequence lengths, for the PE baselines on the PubMed dataset we directly compare against the Aug-PE results from Xie et al. (2024).
@@ C.2 DP-FTRL
[…]
@@ D.1 Privacy Accounting
[…]
@@ D.2.1 LargeFedBench Evaluation
D.2.2 PubMed Evaluation
For PubMed, our Private Evolution baseline compares to Aug-PE, which has already been evaluated on PubMed (Xie et al., 2024). Note that PubMed was used by Xie et al. (2024) to evaluate central DP algorithms. In the central DP setting, there are no clients; all private data is held at the server and the goal is to release a model with DP guarantees. The notion of neighboring dataset in central DP is a centrally held dataset that is the same except for a single data sample. To compare our algorithm directly with results reported for Private Evolution (Aug-PE) (Xie et al., 2024), we replicate the central DP setting for this dataset by having one PubMed abstract per client and sampling all clients every iteration (or “round”, in our case). We do not compare directly with the results reported in the PrE-Text paper (Hou et al., 2024) because they did not release the precise datasets used.

### v2

@@ 5.1 Main Results
Table 1 lists the accuracy (next token prediction and text classification) achieved by baseline methods (DP-FedAvg, DP-FTRL, Private Evolution) and POPri. In this table, we assume full participation (no client sampling) for fair comparison to baselines, some of which do not have client sampling versions. We find that POPri outperforms all the baseline algorithms. Furthermore, in the $\epsilon=1$ setting POPri closes the gap between fully private learning ($\epsilon=0$) and fully non-private learning ($\epsilon=\infty$) by 40-58% depending on the setting, compared to PE which closes 1-28%. For all methods tested, the measured accuracy values do not depend strongly on $\epsilon$. This has been observed in prior work on DP synthetic data using LLMs (Xie et al., 2024; Hou et al., 2024). POPri outperforms Private Evolution (Aug-PE) even when holding our synthetic sample budget to 2000. Note that synthetic samples are cheap in POPri (we could generate many more) because we have access to the full model, while Xie et al. (2024) only assume access to a model API.
@@
For FedAvg, each round the sampled clients download and upload the downstream model, which in our case is DistilGPT2. This is an 82M (82 million) parameter model leading to a download and upload cost of 82M floats. The client runtime cost comes from local gradient computation, and server runtime is negligible because the server only needs to average model deltas from the clients. For PE, the communication cost comes from each client downloading $K=1800$ sentence embeddings of size 384 resulting in a download cost of 700K (700,000) floats, and uploading a histogram of size 1800 resulting in an upload cost of 1800 floats. The client runtime cost comes from calculating a nearest neighbors histogram and the server runtime cost for PE comes mainly from using the LLM $\Psi$ to generate synthetic samples each round. In POPri each client downloads $K\times J=1800\times 10$ sentence embeddings for a download cost of 7M floats and uploads a vector of size 18,000 for an upload cost of 18,000 floats. The client runtime cost of POPri comes from calculating the cosine similarities, and the server runtime comes from both using $\Psi$ to generate synthetic samples and running DPO.
@@ 5.2 Ablations
In each round a fixed number of clients is subsampled uniformly at random for feedback generation. Figure 3 shows the next-token prediction accuracy ($\%$) of four algorithms for different numbers of clients per round. POPri consistently outperforms all of the baselines, regardless of the client sampling rate. Moreover, POPri’s accuracy is not sensitive to the client sampling rate.
@@
Synthetic datasets are often generated using a language model distinct from the one being aligned (Guo et al., 2024), making the alignment phase inherently off-policy as the model evolves during training. This is reflected in the synthetic data, where the FID score (relative to a held-out evaluation set) worsens after improving. Figure 4 shows PCA visualizations of synthetic data embeddings across alignment iterations, while the left panels plot the FID score and median distance to the medoid in the PCA space. The data distribution transitions from being initially clustered to (roughly) matching the true data distribution, back to being clustered, likely due to overfitting. Early stopping based on validation metrics can help.
@@ 6 Conclusion
Private on-device learning is important when data is stored on edge devices with hardware, storage, and privacy constraints. We propose POPri, which recasts synthetic data-based approaches for private learning as an LLM policy optimization problem. POPri makes several novel design choices in how it gathers and utilizes client feedback to generate DP synthetic data, which is used to finetune a downstream on-device model. POPri outperforms DP-FL and synthetic data baselines on a variety of tasks, including on a large-scale LargeFedBench, a new federated benchmark we have curated.
@@
In this paper, we train models satisfying differential privacy guarantees. When using differential privacy as a tool for protecting user data, it is important to communicate to users what the privacy guarantees mean to be able to obtain informed consent. The algorithms in this paper also use LLMs, which were trained on large scale public text data. While this data was public, explicit consent may not have been given for its use in training the models. The algorithms using LLMs in the paper make no claims about the privacy guarantees of data used in the pretraining of the LLMs.
While our work aims to show how synthetic data can be useful for federated learning, it also poses a number of ethical risks, including the generation of biased or harmful content. In particular, our method (and all variants of private evolution) inherits the biases and undesirable aspects of the public LLM. For example, suppose the public LLM only generates text in English, but some clients’ private data is all in Spanish. In these settings, clients would be forced to vote on synthetic samples, even if potentially none of them are relevant to the client. This may cause the client to contribute data reinforcing a model that is actively not useful (or even harmful) to the client. In contrast, DP-SGD methods do not suffer from this shortcoming, because they do not rely on a public LLM. This problem raises an important point—how can we design DP synthetic data algorithms in which clients can stem the biases or failures of the public LLM, based on their own data? This important question is beyond the scope of the current paper.
Acknowledgments
This work was supported in part by NSF grants CCF-2338772 and CNS-2148359, as well as C3.ai, Bosch, Intel, and the Sloan Foundation. This work used Bridges-2 GPU (Brown et al., 2021; Buitrago & Nystrom, 2021) at the Pittsburgh Supercomputing Center through allocation CIS240135 and CIS240937 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296 (Boerner et al., 2023). The authors acknowledge the National Artificial Intelligence Research Resource (NAIRR) Pilot, the AI and Big Data group at the Pittsburgh Supercomputing Center, and NCSA Delta GPU for contributing to this research result.
@@ Appendix A Algorithmic Details
Here we provide the pseudocode for the centralized version of POPri and the central + conditional generation version of POPri.
Algorithm 2 POPri (central DP, unconditional)
1:  Input: Number of iterations $T$, Noise multiplier $\sigma$, LLM $\Psi$, embedding model $\Gamma$, base prompt $\eta$, random prompt generator $\Lambda(\cdot)$, “rejected” index $\ell$, private dataset $S$, $K$ number of prompts, $J$ number of responses per prompt
2:  Output: LLM for generating synthetic data $\Psi_{T+1}$
3:
4:  Embed all private samples $E=\Gamma(S)$
5:  Initialize LLM $\Psi_{1}=\Psi$
6:  for $t\leftarrow 1$ ... $T$  do
7:  Initialize the response vector $R=\emptyset$
8:  for $k\leftarrow 1\ldots K$  do
9:   Generate prompt $\eta_{k}=\Lambda(\eta)$,
10:   Generate $J$ responses $R_{kj}=\Psi_{t}(\eta_{k})$, $j\in[J]$
11:  end for
12:  Calculate embeddings $E_{syn,t}=\{\Gamma(R_{kj})\}_{k\in[K],j\in[J]}$
13:  ${\rm Scores}_{t}\leftarrow$ ${\tt CENTRALSCORE}(E_{syn,t},E)+\mathcal{N}(0,\sigma^{2}{I})$
14:  Set $P[k,j]$ as the $j$-th highest score response for prompt $\eta_{k}$, according to ${\rm Scores}_{t}$
15:  Initialize preference dataset $\mathcal{P}_{t}=\emptyset$
16:  for $k\leftarrow 1\ldots K$  do
17:   Select positive synthetic sample: $\mathcal{P}_{t}[k,1]=P_{t}[k,1]$
18:   Select negative synthetic sample: $\mathcal{P}_{t}[k,2]=P_{t}[k,\ell]$
19:  end for
20:  Fine-tune: $\Psi_{t+1}\leftarrow\text{DPO}(\Psi_{t},\{\eta_{k}\}_{k\in[K]},\mathcal{P}_{t})$
21:  end for
22:  Output $\Psi_{T+1}$
Algorithm 3 POPri (central DP, conditional)
1:  Input: Number of iterations $T$, Noise multiplier $\sigma$, LLM $\Psi$, embedding model $\Gamma$, base prompt $\eta$, conditional (class-specified) random prompt generator $\Lambda(\cdot,\cdot)$, “rejected” index $\ell$, private dataset $S$, $K$ number of prompts, $J$ number of responses per prompt, number of classes $B$
2:  Output: LLM for generating synthetic data $\Psi_{T+1}$
3:  Embed private samples for each class i=1Bi=1...B, $E_{i}=\Gamma(\{s\}_{F(s)=i,s\in S})$ where $F(s)$ is the class index of sample s
4:  Initialize LLM $\Psi_{1}=\Psi$
5:  for $t\leftarrow 1$ ... $T$  do
6:  Initialize B response vectors $R=\{\emptyset,...\emptyset\}=\{R^{(1)},...,R^{(B)}\}$
7:  for $b\leftarrow$ ... $B$ do
8:   for $k\leftarrow 1\ldots K$  do
9:     Generate prompt $\eta_{k}=\Lambda(\eta,b)$,
10:     Generate $J$ responses $R_{kj}^{(b)}=\Psi_{t}(\eta_{k})$, $j\in[J]$
11:   end for
12:   Calculate embeddings $E_{syn,t}^{(b)}=\{\Gamma(R_{kj}^{(b)})\}_{k\in[K],j\in[J]}$
13:   ${\rm Scores}_{t}\leftarrow$ ${\tt CENTRALSCORE}(E_{syn,t}^{(b)},E_{b})+\mathcal{N}(0,\sigma^{2}{I})$
14:   Set $P[k,j]$ as the $j$-th highest score response for prompt $\eta_{k}$, according to ${\rm Scores}_{t}$
15:   Initialize preference dataset $\mathcal{P}_{t}^{(b)}=\emptyset$
16:   for $k\leftarrow 1\ldots K$  do
17:     Select positive synthetic sample: $\mathcal{P}_{t}^{(b)}[k,1]=P_{t}[k,1]$
18:     Select negative synthetic sample: $\mathcal{P}_{t}^{(b)}[k,2]=P_{t}[k,\ell]$
19:     Set prompt $\mathcal{P}_{t}^{(b)}[k,3]=\eta_{k}$
20:   end for
21:  end for
22:  Fine-tune: $\mathcal{P}_{t}=\bigcup_{b=1}^{B}P_{t}^{(b)}$, $\Psi_{t+1}\leftarrow\text{DPO}(\Psi_{t},\mathcal{P}_{t})$
23:  end for
24:  Output $\Psi_{T+1}$
Below is the similarity scoring function we use for the federated setting.
@@
1:  Input: Set of embeddings of private client data $E_{i}=\{emb(s_{1}^{(i)}),\ldots,emb(s_{m_{i}}^{(i)})\}$ for $i\in\mathcal{S}^{t}$}, embeddings of synthetic data ${\rm E}_{syn}$, total synthetic samples $M=K\times J$ $\rm{Scores}$ $\leftarrow\boldsymbol{0}^{M}$
2:  ${\rm Scores}[j]$ = $\frac{1}{m_{i}}\sum_{e_{pri}\in E_{i}}\frac{\langle e_{pri},e_{j}\rangle}{\|e_{pri}\|\|e_{j}\|}$ for $e_{j}\in E_{syn}$
3:  $\bf return$} $\mathbf{\text{Scores}}/{\max\!\bigl{(}1,\;\lVert\mathbf{\text{Scores}}\rVert_{2}\bigr{)}}$
Below is the similarity scoring function we use for the central DP setting.
Algorithm 5 CENTRALSCORE
1:  Input: Embeddings of private data $E$, embeddings of synthetic data ${E}_{syn}$ $\rm{Scores}$ $\leftarrow\boldsymbol{0}^{M}$
2:  ${\rm Scores}[j]$ = $(1/|E|)\sum_{e_{pri}\in E}\frac{\langle e_{pri},e_{j}\rangle}{\|e_{pri}\|\|e_{j}\|}$ for $e_{j}\in E_{syn}$
@@ B.1 Model and Hyperparameters
In each round we generate 18000 synthetic data samples for the clients to evaluate. This is accomplished with 1800 prompts, each generating 10 samples for clients to rank. We select the 1st and 5th ranked sample for a given prompt for the “selected” and “rejected” data samples in the DPO preference dataset. We describe the experiments regarding which rank to use for constructing the preference dataset in detail in Appendix Section E.3. To test the scaling relation with the number of clients per round and the total number of clients participating in the training, we set up the parameters and privacy budget shown in Table D.1. The ‘all-MiniLM-L6-v2’ sentence transformer model is used as the embedding model in POPri. We note that we adopt “sentence-t5-base” sentence transformer for PubMed during the step of fine-tuning ${\rm BERT}_{small}$, which follows the setting in AUG-PE. We ensure POPri follows privacy guarantee of $(\epsilon,\delta)$-DP = (1, $3\times 10^{-6}$) or (7, $3\times 10^{-6}$) for both the bioRxiv and the Congressional Speeches datasets and run with 20 iterations for DP-FedAvg, DP-FTRL, PrE-Text for comparison. For AUG-PE, we set $(\epsilon,\delta)$-DP = (1, $2.72\times 10^{-6}$) or (4, $2.72\times 10^{-6}$). PubMed experiments are run with 10 iterations.
@@
  • For BioRxiv &\& Congressional Speeches, we fine-tuned the pre-trained DistillGPT2 for next-token prediction. We set the max sequence length as 64, number of generated synthetic data as 1,000,000, the batch size as 160, the learning rate as $2e^{-4}$, and the number of epochs as 80.
  • For PubMed, to compare with (Yue et al., 2023b), we follow their procedure to leverage pre-trained ${\rm BERT}_{small}$ (Turc et al., 2019). We set the max sequence length as 512, number of generated synthetic data as 2000, batch size as 32, learning rate as 3e-4, the weight decay as 0.01, and the number of epochs as 10. To compare with (Xie et al., 2024), we set up the $(\epsilon,\delta)$-DP value and hypterparameter according to their choice. For example, they set $\delta={1\over N_{priv}\cdot{\rm log}(N_{priv})}$ following (Yue et al., 2023b). To achieve $\delta$ = $\{$1,4}\}, we use noise multiplier $\sigma$ = $\{$13.7, 3.87}\} for 10 iterations under DP on all PubMed data. Note that our noise multiplier values are slightly different than (Xie et al., 2024) due to different methods for calculating differential privacy.
@@ B.2 Prompt Design
To compare with other data generator methods, we adopt the prompts used in the baseline models against which we compare. We generate the synthetic data using an approach similar to that in PrE-Text (Hou et al., 2024). Figure 5 shows an example of the prompt we use for prompting LLaMA-3-BB for generating synthetic data. For bioRxiv/Congress, we randomly take text samples from the c4 (Raffel et al., 2019) dataset as our examples in the prompt. For PubMed, while running POPri, we still adopt the prompt shown in Figure 5 but reduce the number of examples to two in order to accommodate longer sequence lengths, randomly sampling generated abstracts from LLaMA-3-8B. For OpenReview, we prompt the model directly to generate paper reviews (similarly to (Xie et al., 2024)).
@@ Appendix C Implementation Details of Baseline Models
In this section we provide implementation details for the baseline algorithms. We use two DP-FL baselines: DP-FedAvg and DP-FTRL. For the PE baseline, we implement PrE-Text (Hou et al., 2024) for the evaluations on the bioRxiv and Congressional Speeches datasets. For the PE baselines on the PubMed dataset we directly compare against the Aug-PE results from Xie et al. (2024).
@@ C.2 DP-FTRL
We also use the DP variant of Follow-The-Regularized-Leader (DP-FTRL) algorithm (Kairouz et al., 2021a) to fully fine-tune DistilGPT2. The hyperparameter settings are similar to DP-FedAvg other than the noise multipliers. The noise multipliers are $\sigma$ = $\{$19.5, 3.35}\} when considering all the data, and the settings for partial participation experiments are shown in Table D.1.
@@ D.1 Privacy Accounting
ccccccc Total #\# of #\# of clients $\sigma_{1}\tabularnoteFor$ DP-FedAvg, PrE-Text, POPri., $\epsilon$ = 7 $\sigma_{1}\tabularnoteFor$ DP-FedAvg, PrE-Text, POPri., $\epsilon$ = 1 $\sigma_{2}\tabularnoteFor$ DP-FTRL, $\epsilon$ = 7 $\sigma_{2}\tabularnoteFor$ DP-FTRL, $\epsilon$ = 1
@@ D.2.1 LargeFedBench Evaluation
D.2.2 PubMed and OpenReview Evaluation
For PubMed and OpenReview, our Private Evolution baseline compares to Aug-PE, which has already been evaluated on PubMed and OpenReview (Xie et al., 2024). Note that PubMed and OpenReview was used by Xie et al. (2024) to evaluate central DP algorithms. In the central DP setting, there are no clients; all private data is held at the server and the goal is to release a model with DP guarantees. The notion of neighboring dataset in central DP is a centrally held dataset that is the same except for a single data sample. To compare our algorithm directly with results reported for Private Evolution (Aug-PE) (Xie et al., 2024), we replicate the central DP setting for this dataset by having one PubMed abstract per client and sampling all clients every iteration (or “round”, in our case).